OpenStudy (anonymous):

5a√5a^3b + 2a^2 √45ab − 4 √125a^5b

OpenStudy (neer2890):

is this your question: \[5a \sqrt{5}a ^{3b}+2a ^{2}\sqrt{45}ab-4\sqrt{125}a ^{5}b\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

actually no...the 5(a^3)b

OpenStudy (neer2890):

if this is your question then it wil become \[-4\sqrt{125}a ^{5}b+5\sqrt{5}a ^{4}b+2\sqrt{45}a^{3}b\] now take \[a^{3}b\] common \[a ^{3}b(-4\sqrt{125}a ^{2}+5\sqrt{5}a+2\sqrt{45})\]

OpenStudy (neer2890):

now this will become quadratic equation in a solve this equation and you will get your answer.

OpenStudy (anonymous):

thank you

OpenStudy (neer2890):

you are welcome @maria_valentine11